Circular No. 6543

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997T IN UGC 6896 J. Mueller reports her discovery of an apparent supernova (mag ∼ 17.5)

on a plate taken on Jan. 19 with the 1.2-m Oschin Schmidt telescope in the course of the second Palomar Sky Survey. Mueller reports the following approximate position for the supernova: α = 11^{h}55^{m}48^{s}.8, δ = +80^{o}12′44′′ (equinox 2000.0); this is 9′′ east and 29′′ south of the southern bright component of this three-galaxy complex.

S. Nakano, Sumoto, Japan, reports the following precise position for SN 1997T, measured by Y. Kushida from an unfiltered CCD frame taken on Jan. 22.619 UT by R. Kushida (Yatsugatake South Base Observatory, 0.40-m reflector): α = 11^{h}55^{m}47^{s}.46, δ = +80^{o}12′42′′.6 (equinox 2000.0), mag 17.6; this corresponds to 9′′.5 east and 27′′.3 south of the brightest galaxy component.

XTE J1709–267 F. E. Marshall, Laboratory for High Energy Astrophysics, Goddard

Space Flight Center (LHEA/GSFC); J. H. Swank, LHEA/GSFC; B. Thomas, Hughes STX and LHEA/GSFC; L. Angelini, Universities Space Research Association (USRA) and LHEA/GSFC; A. Valinia, National Research Council and LHEA/GSFC; and K. Ebisawa, USRA and LHEA/ GSFC, report the discovery of an x-ray nova with the Proportional Counter Array (PCA) onboard the Rossi X-ray Timing Explorer (RXTE). The source was first seen for about 20 s on Jan. 16.016 UT during a satellite maneuver and seen again during a second maneuver on Jan. 18.147. The best-fit position for the source is α = 17^{h}09^{m}.3, δ = −26^{o}44′ (equinox 2000.0; estimated systematic uncertainty ∼ 6′). During both observations, the 2–10-keV source intensity was ∼ 0.14 times that of the Crab. There is no catalogued hard x-ray source near this position, but the source may be associated with the ROSAT All-Sky Survey (IAUC 6420) source 1RXS J170930.2–263927, which is 5′ from the best-fit RXTE position. The ROSAT source was 13 mCrab in the band 1–2.4 keV. Column densities of spectral fits to the PCA slew data depend on the model; a power law has a photon index of 3.06 ± 0.06 and absorption by a column density of (5.8 ± 0.6) × 10^{22} cm^{−2}, while a bremsstrahlung model has kT = 4.7 ± 0.2 keV and half the column density. Additional observations are encouraged, to determine an accurate position of the new source, test the possible association with J170930.2–263927, and determine other properties.”

1997 January 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green